I do not understand how my instructor got the answer to this question of a logarithmic equation. log(5x-4)divided by logx =2 . The answer is 4. When I take 10 as the base of the log I get 5x-4=100x and this does not make sense. Can anyone help?
Answer by richwmiller(17219)

base 10
What is it divided by? (logx)-2 or log (x-2)
if it is (log(x))-2 the result is 4
log(5x-4)/logx -2=0
add 2 to both sides
log(5x-4)/logx=2
multiply both sides by logx
log(5x-4)=2(logx)
raise both to power of e
5x-4=x^2
solve the quadratic
solution 1 doesn't work
solution 4 does